OpenStudy (anonymous):

Solve 4(5m + 6) - 8 <= 336

OpenStudy (tkhunny):

Have you considered adding 8 and then dividing by 4?

OpenStudy (pinksapphire):

m<=16

OpenStudy (tkhunny):

@PinkSapphire Please read the Code of Conduct and heed its wisdom. Right or wrong, that is not an appropriate response.

OpenStudy (pinksapphire):

What?

OpenStudy (tkhunny):

http://openstudy.com/code-of-conduct

OpenStudy (pinksapphire):

Oh okay, thank you...I won't do that again :)

OpenStudy (tkhunny):

Lot's of folks can use a willing helper. Keep contributing!

OpenStudy (pinksapphire):

Should I show my work?

OpenStudy (tkhunny):

Excellent question! Always. Provide explanations and thinking. We need to teach understanding, not just the art of checking the right box.

OpenStudy (pinksapphire):

4(5m+6)-8<=336 (20m+24)-8<=336 20m+16<=336 -16 -16 20m<=320 20m/20<=320/20 m<=16

OpenStudy (tkhunny):

I would have done it a little differently. My way isn't "better", per se, it just has a little different emphasis. My rule of thumb is to make the numbers we start with the worst ones we see. 4(5m+6)-8<=336 Add 8 4(5m+6)<=344 Divide by 4 5m+6 <= 86 Subtract 6 5m <= 80 Divide by 5 m <= 16 Unique answers don't care how you find them! It is a beautiful thing.